# Anomalous enhancement of Néel order in the $S=1/2$ square lattice Heisenberg model under fictitious magnetic field

### v1

@@ III.1 Spin-wave approximation
where $m_{s}(0)=0.303$. As shown in Fig. 4(b), the difference $\Delta m_{s}$ behaves as $\Delta m_{s}\propto\sqrt{\phi}=l_{B}^{-1}$ for small magnetic fluxes. Generally, a physical quantity could be expanded in integer powers of $l_{B}^{-1}$ (with possible logarithmic corrections) within the spin-wave approximation which is a free bosons description, since the $l_{B}$ is the cut off scale for the gapped magnons as mentioned above. The $\Delta m_{s}\propto l_{B}^{-1}$ dependence is the leading order behavior with respect to $l_{B}$. In this sense, $\Delta m_{s}$ can be regarded as a correction to $m_{s}(0)$ due to the finite magnetic length $l_{B}$ just like the conventional finite size effect for a system with a linear size $L$ [46]. In Appendix B, we extend our discussion to the XXZ model with Ising anisotropy and demonstrate that the system is indeed governed by the magnetic length $l_{B}$ based on a finite-size scaling analysis. Note that $\Delta E$ and $\Delta m_{s}$ slightly deviate from the exact power-law with the exponent $1/2$ in the very small region of $0\leqq\phi\lesssim 0.01$. Two possible reasons for this deviation can be considered. The first one is that the magnetic length $l_{B}$ may exceed the system size $L$, potentially leading to non-critical behaviors in the weak magnetic field region. The second is a technical issue regarding instability in the spin-wave calculation due to specific modes which are connected to the gapless modes at $\phi=0$. The deviation is not physically meaningful, because these problems can be avoided in the thermodynamic limit.
From the spin-wave calculation of the $z$-axis Néel order, one can see that the fictitious magnetic field effectively enhances the Ising interaction $JS^{z}_{i}S^{z}_{j}$. One can compare the results above for the Heisenberg model at $\phi\neq 0$ with the previous results for the XXZ model at $\phi=0$, since these two models have the same symmetry ${\mathrm{U}}(1)\times{\mathbb{Z}}_{2}$. The XY interaction is denoted by $J_{xy}$ and the Ising interaction by $J_{z}\geq J_{xy}$. In the XXZ model at $\phi=0$, the spin-wave approximation gives $\Delta E(J_{z})\propto\sqrt{J_{z}-J_{xy}}$ and $\Delta m_{s}(J_{z})\propto\sqrt{J_{z}-J_{xy}}$ in the leading order [47, 48]. The $\sqrt{J_{z}-J_{xy}}$ dependence in these quantities has been confirmed by various methods [44, 45], and one would naively expect that the $\sqrt{\phi}$ dependence in the Heisenberg model at $\phi\neq 0$ holds beyond the spin-wave approximation as well. However, magnon interactions are neglected in the spin-wave approximation and validity of this approximation may be non-trivial in the present system. To resolve this problem, we will discuss the DMRG calculations in the next section.
Figure 5: $(m^{0}_{\rm corr})^{2}$ plotted versus $1/L$ for $L=4,6,8,10$. The dashed curve is a polynomial fitting, $m=m_{0}+m_{1}/L+m_{2}/L^{2}+m_{3}/L^{3}$, which gives $\left(m_{0}\right)^{2}=0.087(1)$ and $\left(m_{0}\right)=0.295(3)$.
@@ III.2 DMRG calculation
Figure 6: Néel correlation function $\left(m^{z}_{\rm corr}\right)^{2}$ as a function of $\phi$ for various system sizes $(L_{x},L_{y})=(2L,L)$ calculated by DMRG. The yellow square is for $L=6$, and the green triangle is for $L=8$, and the orange cross is for $L=10$.
We find that the calculated ground state at $\phi\neq 0$ is essentially a uniform state with the long range Néel order similarly to the results within the spin-wave approximation in the previous section, although there are some deviations around the open boundaries. The ground state belongs to the $S^{z}_{\rm tot}=\sum_{i}S^{z}_{i}=0$ sector.
Figure 6 shows the calculated $(m^{z}_{\rm corr})^{2}$ as a function of the fictitious magnetic flux $\phi$. Similarly to the results by the spin-wave approximation in the previous section, the Néel magnetization is enhanced by the flux. Besides, the system size dependence becomes smaller as $\phi$ increases, which is qualitatively consistent with the finite correlation length $\sim l_{B}$ of the gapped magnons in the spin-wave approximation since finite size effects are generally suppressed in a gapped system. However, the two results obtained by the spin-wave approximation and the DMRG are significantly different in a quantitative manner. Namely, the enhancement of the Néel magnetization around $\phi=0$ in the DMRG calculation is much stronger than that in the spin-wave approximation. This behavior could be described by a power-law function with a small exponent or by a discontinuous function.
@@
where $m^{z}_{\rm corr}(0)=m^{0}_{\rm corr}(0)/\sqrt{3}\equiv 0.295/\sqrt{3}$. Note that $\Delta m^{z}_{\rm corr}(0)=0$ in the thermodynamic limit by definition, (See also Appendix C.) We show a logarithmic scale plot of the difference $\Delta m^{z}_{\rm corr}$ in Fig. 7. It has only small $L$-dependence for relatively large magnetic fields as already mentioned. In this field region, $\Delta m^{z}_{\rm corr}$ exhibits a scaling behavior,
$\Delta m^{z}_{\rm corr}=C\cdot\phi^{n},$ (7)
where the prefactor $C$ only weakly depends on the system size $L$. This scaling behavior holds above a crossover scale $\phi_{L}$ for each system size, and it is roughly estimated to be $\phi_{6}\simeq 0.05,\phi_{8}\simeq 0.04,\phi_{10}\simeq 0.03$.
@@
We obtain the exponent $n$ and the prefactor $C$ for each system size $L$, and they are extrapolated for $L\to\infty$ (Fig. 8). These results are summarized in Table 1. Numerically, the crossover scale $\phi_{L}$ decreases as the system size increases, $\phi_{6}>\phi_{8}>\phi_{10}$. Although we naively expect that $\phi_{L}\to 0$ as $L\to\infty$, it could converge to some non-zero value $\phi_{\infty}$ and a different scaling exponent might be obtained for $\phi<\phi_{\infty}$. In any case, the obtained exponent $n\simeq 0.136$ for $L=6,8,10$ gives an upper bound since the slope in Fig. 7 for $\phi<\phi_{0}$ is smaller than that for $\phi>\phi_{0}$. Therefore, we conclude that the fictitious magnetic field leads to the anomalously strong enhancement of the Néel order with the scaling behavior $\Delta m^{z}_{\rm corr}\propto\phi^{n}$ with $n\lesssim 0.136$. Note that the above discussion is based on the assumption that the enhancement of $m^{z}_{\rm corr}(\phi)$ is continuous and can be described by a power-law function. If $m^{z}_{\rm corr}(\phi)$ is discontinuous at $\phi=0$, the enhancement is even stronger than any power-law behavior. Since the Néel order is continuous within the spin-wave approximation as demonstrated in the previous section, a possible discontinuity will be understood as a fluctuation-induced discontinuity. In any case, our DMRG calculation suggests importance of the magnon interactions neglected in the spin-wave approximation. This may be somewhat counter-intuitive, because magnon interactions are usually negligible at low temperatures where there are few thermally excited magnons. Indeed, the significant difference between the spin-wave results and the DMRG results is highly non-trivial even when compared with the magnetic catalysis in any known systems, as discussed in the next section.
Figure 8: The exponent $n$ plotted versus $1/L$ with $L=6,8,10$. The dashed line represents the linear fitting with weighted least squares, which gives $n\simeq 0.136$ in the thermodynamic limit.
@@
Table 1: Results of the power-law fitting Eq. (7). The exponent $n$ in the thermodynamic limit is estimated with weighted least squares. The prefactor $C$ is estimated to be nearly 0.286 in the thermodynamic limit, as the finite size effect is negligible within the present numerical accuracy.
@@ III.3 Discussion based on magnetic catalysis
Here, we give a brief alternative discussion for the anomalous enhancement of the Néel order in the Heisenberg model under the fictitious magnetic fields. We will argue that it is beyond the well known magnetic catalysis for correlated systems under magnetic fields [28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. Our discussion is based on the hard-core bosons which are equivalent to $S=1/2$ spins via the relations $c_{j}=S_{j}^{-}$ and $n_{j}=c_{j}^{\dagger}c_{j}=S_{j}^{z}+1/2$, where $c_{j}$ is the boson annihilation operator. The Heisenberg model at $S^{z}_{\rm tot}=0$ can be rewritten as the extended Hubbard model of hard-core bosons at half-filling,
@@
with $2t=V=J$. The hard-core condition ($n_{j}=0,1$) can be regarded as the strong coupling limit of the on-site Hubbard interaction $Un_{j}(n_{j}-1)$. The boson model has the particle number U(1) symmetry and the anti-unitary particle-hole symmetry $z\to z^{*},c_{j}\to c_{j}^{{\dagger}},n_{j}\to 1-n_{j}$ corresponding to the $\mathrm{U}(1)\times{\mathbb{Z}}_{2}$ spin rotation symmetry of the Heisenberg model. The particle-hole symmetry is spontaneously broken in the charge-density wave state as in the $z$-axis Néel state of the Heisenberg model.
@@
For the hard-core boson model, the fictitious magnetic field is nothing but a standard magnetic field for charged bosons or may be regarded as an artificial field in cold atoms [49, 50, 51]. The magnetic field leads to the Landau quantization of single-particle energies and the band structure becomes flat, which means that the hopping integral $t$ is effectively reduced. Consequently the interaction $Vn_{i}n_{j}$ can be enhanced compared to the kinetic energy. To be more precise, the particles are localized within a length scale $\sim l_{B}$, and the system size effectively becomes $\sim l_{B}\times l_{B}$ in two dimensions, where the correlation effects are amplified due to the spatial confinement. Such a mechanism for enhanced correlation effects by magnetic fields is called magnetic catalysis and it has been extensively studied mainly for fermion systems [28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. The magnetic catalysis has been examined also for bosons at finite temperature [52, 53]. In this picture, it is natural to have an enhanced charge density wave order of the hard-core bosons for the Hamiltonian (8). However, there is a striking difference between the magnetic catalysis for the present system and the known systems in a quantitative manner.
One can compare the DMRG results of the Heisenberg model which is located at a bicritical fixed point of the renormalization group [54] with the known magnetic catalysis for interacting fermion systems at critical fixed points. In the $(2+1)$-dimensional Dirac fermions, the magnetic catalysis of an order parameter $\Delta m$ at the semimetal-insulator transition point behaves as $\Delta m\propto l_{B}^{-\beta/\nu}=\phi^{\beta/2\nu}$ [37], and it is described by the exactly same scaling form $\Delta m(L)\propto L^{-\beta/\nu}$ for a finite size system with the linear size $L$ at $\phi=0$ [55, 56, 57, 58, 59]. The exponents $\beta$ and $\nu$ are defined as $\Delta m\sim g^{\beta}$ and $\xi\sim g^{-\nu}$ with $g$ and $\xi$ the reduced interaction and the correlation length, respectively, when there are no other finite size effects. They are $\beta=\nu=1$ in the mean field approximation, and the order parameter behaves as $\Delta m\sim l_{B}^{-1}=\sqrt{\phi}$. According to the extensive numerical studies, the critical exponents are $(\beta,\nu)\simeq(0.53,0.80)\sim(0.63,0.78)$ for a ${\mathbb{Z}}_{2}$ symmetry breaking order parameter [58, 59], and correspondingly, $\Delta m\sim l_{B}^{-0.81\sim-0.66}=\phi^{0.33\sim 0.40}$ [37]. This means that the quantum fluctuations neglected in the mean field approximations play important roles in the magnetic catalysis especially at a quantum critical point, and they lead to strong enhancement of the order parameter by the magnetic field.
@@
We naively expect that the magnetic length $l_{B}$ is a scaling variable in the present hard-core bosons as well, if the order parameter is continuous at $\phi=0$. (More precisely, we consider the singular part of the ground state energy measured from that at $g=l_{B}^{-1}=0$, $\Delta\varepsilon(h,g,l_{B}^{-1})=\varepsilon(h,g,l_{B}^{-1})-\varepsilon(h,0,0)$, where $h$ is an external field conjugate to the Néel magnetization and $g=(V-V_{c})/V_{c}$.) Indeed, in the spin-wave approximation, the Néel order parameter behaves as $\Delta m_{s}^{z}\sim l_{B}^{-\beta/\nu}=1/l_{B}$ similarly to the well known finite size correction $\Delta m_{s}^{z}\sim 1/L$ [46, 44, 45] (see also Appendix B). This is fully consistent with the physical picture of the critical magnetic catalysis for the hard-core bosons with $\beta=\nu=1/2$. However, the critical behavior of the magnetic catalysis in the DMRG calculation is much stronger, $\Delta m^{z}_{\rm corr}\sim\phi^{0.136}=l_{B}^{-0.272}$, if it is continuous at $\phi=0$. Obviously, the simple correspondence between $l_{B}$ and $L$ does not hold, and the scaling dimension of $\phi$ in $\Delta\epsilon(h,g,\phi)$ is $\lambda_{\phi}\neq-1/2$. This implies that the anomalous enhancement of $\Delta m_{\rm corr}^{z}$ cannot be understood based on the standard magnetic catalysis. This may be related to the fact that the fictitious magnetic field leads not only to the Landau quantization but also to the symmetry reduction from SU(2) to $\mathrm{U}(1)\times\mathbb{Z}_{2}$, which is distinct from the electromagnetic field. The Bosonic nature of the constituent particles would also be important for the anomalous scaling.
@@
Another possibility is that $\Delta m_{\rm corr}^{z}$ has the fluctuation induced discontinuity at $\phi=0$, as already mentioned. In any case, the anomalous enhancement of the Néel order (or the charge density wave order in terms of the bosons) is beyond the known magnetic catalysis. It is an interesting future problem to fully clarify the origin and mechanism of the anomalous behavior of $\Delta m^{z}_{\rm corr}(\phi)$.

### v2

@@ III.1 Spin-wave approximation
where $m_{s}(0)=0.303$. As shown in Fig. 4(b), the difference $\Delta m_{s}$ behaves as $\Delta m_{s}\propto\sqrt{\phi}=l_{B}^{-1}$ for small magnetic fluxes. Generally, a physical quantity could be expanded in integer powers of $l_{B}^{-1}$ (with possible logarithmic corrections) within the spin-wave approximation, which is a free bosons description, since $l_{B}$ is the cut off scale for the gapped magnons, as mentioned above. The $\Delta m_{s}\propto l_{B}^{-1}$ dependence is the leading-order behavior with respect to $l_{B}$. In this sense, $\Delta m_{s}$ can be regarded as a correction to $m_{s}(0)$ due to the finite magnetic length $l_{B}$. just like the conventional finite size effect for a system with a linear size $L$ [47]. In Appendix B, we extend our discussion to the $XXZ$ model with Ising anisotropy and demonstrate that the system is indeed governed by the magnetic length $l_{B}$, based on a finite-size scaling analysis. Note that $\Delta E$ and $\Delta m_{s}$ slightly deviate from the exact power-law with the exponent $\frac{2}{2}$ in the very small region of $0\leqq\phi\lesssim 0.01$. Two possible reasons for this deviation can be considered. The first one is that the magnetic length $l_{B}$ may exceed the system size $L$, potentially leading to noncritical behaviors in the weak magnetic field region. The second is a technical issue regarding instability in the spin-wave calculation due to specific modes which are connected to the gapless modes at $\phi=0$. The deviation is not physically meaningful, because these problems can be avoided in the thermodynamic limit.
From the spin-wave calculation of the $z$-axis Néel order, one can see that the fictitious magnetic field effectively enhances the Ising interaction $JS^{z}_{i}S^{z}_{j}$. One can compare the results above for the Heisenberg model at $\phi\neq 0$ with the previous results for the $XXZ$ model at $\phi=0$, since these two models have the same symmetry ${\mathrm{U}}(1)\times{\mathbb{Z}}_{2}$. The $XY$ interaction is denoted by $J_{xy}$ and the Ising interaction by $J_{z}\geq J_{xy}$. In the $XXZ$ model at $\phi=0$, the spin-wave approximation gives $\Delta E(J_{z})\propto\sqrt{J_{z}-J_{xy}}$ and $\Delta m_{s}(J_{z})\propto\sqrt{J_{z}-J_{xy}}$ in the leading order [48, 49]. The $\sqrt{J_{z}-J_{xy}}$ dependence in these quantities has been confirmed by various methods [45, 46], and one would naively expect that the $\sqrt{\phi}$ dependence in the Heisenberg model at $\phi\neq 0$ holds beyond the spin-wave approximation as well. However, magnon interactions are neglected in the spin-wave approximation, and validity of this approximation may be nontrivial in the present system. To resolve this problem, we will discuss the DMRG calculations in the next section.
[…]
@@ III.2 DMRG calculation
The obtained $m_{\rm corr}^{0}$ is slightly smaller than the previous results $m_{\rm corr}^{0}=0.306-0.308$ evaluated by several numerical methods,  [41, 42, 43, 44, 45, 46], because the system sizes used in our extrapolation are limited to $L=4-10$. Nevertheless, the deviation is at most few percents and we have confirmed that it has only negligible influence in the following discussions for $\phi\neq 0$ (see Appendix C).
Figure 6: Néel correlation function $\left(m^{z}_{\rm corr}\right)^{2}$ as a function of $\phi$ for vs system sizes $(L_{x},L_{y})=(2L,L)$ calculated by DMRG. The green square is for $L=6$, and the red triangle is for $L=8$, the purple star is for $L=10$.
We find that the calculated ground state at $\phi\neq 0$ is essentially a uniform state with the long-range Néel order like the results within the spin-wave approximation in the previous section, although there are some deviations around the open boundaries. The ground state belongs to the $S^{z}_{\rm tot}=\sum_{i}S^{z}_{i}=0$ sector.
Figure 6 shows the calculated $(m^{z}_{\rm corr})^{2}$ as a function of the fictitious magnetic flux $\phi$. Like the results by the spin-wave approximation in the previous section, the Néel magnetization is enhanced by the flux. Additionally, the system size dependence becomes smaller as $\phi$ increases, which is qualitatively consistent with the finite correlation length $\sim l_{B}$ of the gapped magnons in the spin-wave approximation since finite size effects are generally suppressed in a gapped system. However, the two results obtained by the spin-wave approximation and the DMRG are significantly different in a quantitative manner. Namely, the enhancement of the Néel magnetization around $\phi=0$ in the DMRG calculation is much stronger than that in the spin-wave approximation. This behavior could be described by a power-law function with a small exponent or by a discontinuous function.
@@
where $m^{z}_{\rm corr}(0)=m^{0}_{\rm corr}(0)/\sqrt{3}\equiv 0.295/\sqrt{3}$. Note that $\Delta m^{z}_{\rm corr}(0)=0$ in the thermodynamic limit by definition (see also Appendix C). We show a logarithmic scale plot of the difference $\Delta m^{z}_{\rm corr}$ in Fig. 7. It has only small $L$ dependence for relatively large magnetic fields as already mentioned. In this field region, $\Delta m^{z}_{\rm corr}$ exhibits a scaling behavior,
$\Delta m^{z}_{\rm corr}=C\phi^{n},$ (7)
where the prefactor $C$ only weakly depends on the system size $L$. This scaling behavior holds above a crossover scale $\phi_{L}$ for each system size, and it is roughly estimated to be $\phi_{6}\simeq 0.05,\phi_{8}\simeq 0.04,\text{and}\phi_{10}\simeq 0.03$.
@@
We obtain the exponent $n$ and the prefactor $C$ for each system size $L$, and they are extrapolated for $L\to\infty$ (Fig. 8). These results are summarized in Table 1. Numerically, the crossover scale $\phi_{L}$ decreases as the system size increases $\phi_{6}>\phi_{8}>\phi_{10}$. Although we naively expect that $\phi_{L}\to 0$ as $L\to\infty$, it could converge to some nonzero value $\phi_{\infty}$, and a different scaling exponent might be obtained for $\phi<\phi_{\infty}$. In any case, the obtained exponent $n\simeq 0.136$ for $L=6,8,10$ gives an upper bound since the slope in Fig. 7 for $\phi<\phi_{0}$ is smaller than that for $\phi>\phi_{0}$. Therefore, we conclude that the fictitious magnetic field leads to the anomalously strong enhancement of the Néel order with the scaling behavior $\Delta m^{z}_{\rm corr}\propto\phi^{n}$ with $n\lesssim 0.136$. Note that the above discussion is based on the assumption that the enhancement of $m^{z}_{\rm corr}(\phi)$ is continuous and can be described by a power-law function. The successful power-law fitting implies that this is indeed the case, and the $\phi$ dependence in our system is distinguished from the $J_{z}$ dependence of the $XXZ$ model. If $m^{z}_{\rm corr}(\phi)$ is discontinuous at $\phi=0$, the enhancement is even stronger than any power-law behavior. Such possible discontinuous behavior may be understood as a fluctuation-induced discontinuity. In any case, our DMRG calculation suggests the importance of the magnon interactions neglected in the spin-wave approximation. This may be somewhat counterintuitive, because magnon interactions are usually negligible at low temperatures where there are few thermally excited magnons. Indeed, the significant difference between the spin-wave results and the DMRG results is highly nontrivial even when compared with the magnetic catalysis in any known systems, as discussed in the next section.
Figure 8: The exponent $n$ plotted vs $1/L$ with $L=6,8,\text{and}10$. The dashed line represents the linear fitting with weighted least squares, which gives $n\simeq 0.136$ in the thermodynamic limit.
@@
Table 1: Results of the power-law fitting Eq. (7). The exponent $n$ in the thermodynamic limit is estimated with weighted least squares. The prefactor $C$ is estimated to be $\sim 0.286$ in the thermodynamic limit, as the finite-size effect is negligible within the present numerical accuracy.
@@ III.3 Discussion based on magnetic catalysis
Here, we give a brief alternative discussion for the anomalous enhancement of the Néel order in the Heisenberg model under the fictitious magnetic fields. We will argue that it is beyond well-known magnetic catalysis for correlated systems under magnetic fields, although the spin-wave results can be understood based on this mechanism [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. Our discussion is based on the hard-core bosons which are equivalent to $S=\frac{1}{2}$ spins via the relations $c_{j}=S_{j}^{-}$ and $n_{j}=c_{j}^{\dagger}c_{j}=S_{j}^{z}+\frac{1}{2}$, where $c_{j}$ is the boson annihilation operator. The Heisenberg model at $S^{z}_{\rm tot}=0$ can be rewritten as the extended Hubbard model of hard-core bosons at half-filling,
@@
with $2t=V=J$. The hard-core condition ($n_{j}=0,1$) can be regarded as the strong coupling limit of the on-site Hubbard interaction $Un_{j}(n_{j}-1)$. The boson model has particle number U(1) symmetry and anti unitary particle-hole symmetry $z\to z^{*},c_{j}\to c_{j}^{{\dagger}},\text{and}\,n_{j}\to 1-n_{j}$ corresponding to the $\mathrm{U}(1)\times{\mathbb{Z}}_{2}$ spin-rotation symmetry of the Heisenberg model. The particle-hole symmetry is spontaneously broken in the charge-density wave state as in the $z$-axis Néel state of the Heisenberg model.
@@
For the hard-core boson model, the fictitious magnetic field is nothing but a standard magnetic field for charged bosons or may be regarded as an artificial field in cold atoms [38, 39, 50]. The magnetic field leads to the Landau quantization of single-particle energies and the band structure becomes flat, which means that the hopping integral $t$ is effectively reduced. Consequently, the interaction $Vn_{i}n_{j}$ can be enhanced compared with the kinetic energy. To be more precise, the particles are localized within a length scale $\sim l_{B}$, and the system size effectively becomes $\sim l_{B}\times l_{B}$ in two dimensions, where the correlation effects are amplified due to the spatial confinement. Such a mechanism for enhanced correlation effects by magnetic fields is called magnetic catalysis and it has been extensively studied mainly for fermion systems [27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. Magnetic catalysis has also examined for bosons at finite temperature [51, 52]. In this picture, it is natural to have an enhanced charge density wave order of the hard-core bosons for the Hamiltonian in Eq.(8).
@@
We naively expect that the magnetic length $l_{B}$ is a scaling variable in the present hard-core bosons as well, if the order parameter is continuous at $\phi=0$. Indeed, in the spin-wave approximation, the Néel order parameter behaves as $\Delta m_{s}^{z}\sim l_{B}^{-\beta/\nu}=1/l_{B}$ , like the well-known finite-size correction $\Delta m_{s}^{z}\sim 1/L$ [47, 45, 46] (see also Appendix B). This is fully consistent with the physical picture of the critical magnetic catalysis seen in mean-field approximation for $(2+1)$-dimensional Dirac electron systems [36]. However, the critical behavior of the magnetic catalysis in the DMRG calculation is much stronger, $\Delta m^{z}_{\rm corr}\sim\phi^{0.136}=l_{B}^{-0.272}$, if it is continuous at $\phi=0$. Obviously, the simple correspondence between $l_{B}$ and $L$ does not hold, and this implies that the anomalous enhancement of $\Delta m_{\rm corr}^{z}$ cannot be understood based on the standard magnetic catalysis. This may be related to the fact that the fictitious magnetic field leads not only to the Landau quantization but also to the symmetry reduction from SU(2) to $\mathrm{U}(1)\times\mathbb{Z}_{2}$, which is distinct from the electromagnetic field. The bosonic nature of the constituent particles would also be important for the anomalous scaling.
@@
Another possibility is that $\Delta m_{\rm corr}^{z}$ has fluctuation-induced discontinuity at $\phi=0$, as already mentioned. In any case, the anomalous enhancement of the Néel order (or the charge-density wave order in terms of the bosons) is beyond the known magnetic catalysis. It is an interesting future problem to fully clarify the origin and mechanism of the anomalous behavior of $\Delta m^{z}_{\rm corr}(\phi)$.